OpenStudy (luckycoins888):

fan and medial The attic of a house has the shape of a triangular prism. The triangular end of the attic has a base that is 12 feet across and a height of 8 feet. The length of the attic is 30 feet from front to back. What is the volume of the attic? ft3

OpenStudy (luckycoins888):

@Queenofwarfare1 @wio @eliassaab

OpenStudy (luckycoins888):

@tejasvir

OpenStudy (anonymous):

look at the formula for triangular prism :)

OpenStudy (danjs):

(area of a triangle) * depth

OpenStudy (danjs):

If you ever have a shape that is just extruded, just find the area of the 2-d shape and multiply it by the extrusion length/depth

OpenStudy (luckycoins888):

i still don't get it

OpenStudy (danjs):

What is the area of a triangle?

OpenStudy (danjs):

\[A = \frac{ 1 }{ 2 }*base*height\]

OpenStudy (luckycoins888):

hieght x width?

OpenStudy (danjs):

So the area of the triangle end is, 1/2 * 12 * 8 = 1/2 * 96 = 48 ft^2

OpenStudy (danjs):

So the volume is, V = (area of triangle) x depth of the attic

OpenStudy (danjs):

= 48 ft^2 * 30 ft = 1440 ft^3

OpenStudy (luckycoins888):

thanks so much

OpenStudy (danjs):

np

OpenStudy (anonymous):

48 ft^2 * 30 ft = 1440 ft^3